# Model Restrictiveness in Functional and Structural Settings

### v2

@@ 4 Related Concepts
A key aspect of our framework is to treat the discrepancy $d$ as a design choice, to be selected based on the intended interpretation and the empirical context, rather than inherited from existing complexity measures that were built for different goals. Concretely:
  • In the certainty-equivalent and discrete-choice applications, we use squared $L^{2}$ discrepancies on certainty-equivalent functions or choice probabilities, so that $d(f,g)$ can be read as an average prediction error with clear economic units.
  • In structural models with endogeneity or multiple equilibria, we work with discrepancies between implied conditional distributions (in reduced form or structural forms), again chosen for their interpretability in terms of approximation error rather than for reproducing any particular capacity bound.
Our restrictiveness measure is defined at the level of prediction rules and a user-chosen discrepancy function $d$, with no reference to a particular sample size or estimation procedure. This section discusses and develops formal connections between restrictiveness and related concepts in econometrics, statistics and machine learning.
@@ 4.1 Rademacher Complexity and VC Dimensions
This subsection explains how standard capacity measures from learning theory fit within our framework, and why their asymptotic behavior is driven by the discrepancies they encode.
Following Ellis and Neff (2025), suppose that $\mathcal{X}$ is a closed and bounded infinite subset of $\mathbb{R}^{m}$ for some finite $m\in\mathbb{N}$, $X$ be a random vector of $\mathcal{X}$ with distribution $P_{X}$, and that the outcome space is $\mathcal{Y}=\{-1,1\}$ [^4] Ellis and Neff allow the outcome $Y$ to take values in [1,1][-1,1]. For expositional purposes, we restrict attention to the canonical binary case, since the most standard definitions of both the Rademacher complexity and the VC dimension are formulated for binary-valued function classes.. Let $\bar{\mathcal{F}}_{\sigma}=\{f:\mathcal{X}\to\{-1,1\}\}$ be the set of all deterministic binary classifiers, and let $\mathcal{F}_{\Theta}\subseteq\bar{\mathcal{F}}_{\sigma}$ be a model class (for example, a parametric class indexed by $\theta\in\Theta$). The paper then takes the eligible set to be the full binary class $\mathcal{F}:=\bar{\mathcal{F}}_{\sigma}$ with some probability measure $\lambda_{\mathcal{F}}$; and the discrepancy to be the correlation-based quantity
@@
As we show in Appendix D, since $f$ and $g$ are binary-valued, the discrepancy $d_{\text{VC}}(f,g)=\mathbb{P}(f(X)\neq g(X))$ satisfies $d_{\text{Rad}}=2d_{\text{VC}}$, so it induces exactly the same restrictiveness ordering. Ellis and Neff (2025) establishes that under, this discrepancy, all model classes with finite VC dimension are fully restrictive in the limit. Of course, Rademacher complexity and VC dimension are designed to control worst-case generalization error under adversarial labelings, so it is natural that produce degenerate restrictiveness measures in infinite domains. This degeneracy should thus be interpreted as a consequence of the underlying discrepancy rather than as a limitation of restrictiveness itself. An important point of our paper is that the the discrepancy function $d$ should be chosen to maximize interpretability in a user-driven context-specific manner.
We discuss a related, though slightly different point, about the choice of discrepancy function in the next subsection.
@@ 4.2 Discrepancy Associated with GMM Criterion Function
In many structural econometric models, parameters are estimated by minimizing the sample analog of a moment-based criterion function, such as the generalized method of moments (GMM) criterion. It is natural to ask whether such criteria implicitly define a discrepancy that can be used to evaluate restrictiveness. This subsection shows that  the discrepancy function implicitly associated with the GMM criterion is focused on measuring violations of moment conditions, but not on the distance between prediction rules. Thus, although the GMM criterion is convenient for econometric identification and estimation, the resulting discrepancy is hard to interpret for our purposes.
@@
A natural temptation is to interpret $Q_{\text{GMM}}(\theta)$ as defining a discrepancy between $m_{\theta}$ and the true regression function $g$, and hence to define
@@
unless the instruments are sufficiently rich to identify $g$ pointwise. Thus as shown in the next example $Q_{\text{GMM}}$ does not satisfy $d(f,g)=0\Rightarrow f=g\ \text{(almost surely)}$ which we require of our discrepancies.
@@ 4.3 Limit of the Average-Case Learning Curve
Proposition 2 shows that, under mild assumptions, restrictiveness can be interpreted as the ratio of two long-run average-case function-space approximation errors: one achieved by the best-fitting member of the model class $\mathcal{F}_{\Theta}$, the other by the baseline model class $\mathcal{F}_{\text{base}}$. Low restrictiveness corresponds to a model whose best-fitting members achieve low asymptotic risk across pseudo-truths $f\sim\lambda_{\mathcal{F}}$, while high restrictiveness corresponds to a model that, on average, cannot reduce risk much relative to the baseline. This is analogous to but also different from the standard notion of learning curve in machine learning, which features irreducible errors arising from the noise term $\epsilon_{i}$ in the data generating process $Y_{i}=f(X_{i})+\epsilon_{i}$, which contrasts with our noise-free setup (22). Hence, restrictiveness can be viewed as a normalized noise-free analog of the limit of the standard average-case learning curve. This again illustrates our point that restrictiveness seeks to reveal that functional-form flexibility of the model in question.
See Appendix E for a more detailed explanation of the relationship and differences between restrictiveness and the standard learning curve, in the context of Gaussian process regressions.
@@ 5 Applications
This section applies our restrictiveness framework to three settings: certainty equivalents, multinomial choice with exogenous product characteristics, and multinomial choice with endogenous prices. In the certainty-equivalent application, the relative importance of parameters mirrors that in FGL, but restrictiveness is uniformly higher when models are evaluated over the full continuum of lotteries. In discrete-choice models without endogeneity, restrictiveness is driven primarily by flexibility in mean utility, and the restrictiveness of commonly used empirical specifications differ meaningfully despite the theoretical generality of mixed logit. When prices are endogenous, moment restrictions substantially increase restrictiveness and alter model rankings. Together, the first two applications illustrate how the framework evaluates restrictiveness for functional models on continuum domains, while the third shows how it extends naturally to structural, semiparametric models with endogeneity.
@@ 5.1 Cumulative Prospect Theory
In the first example, we revisit the evaluation of “CPT”, a popular three-parameter specification of Cumulative Prospect Theory (Tversky and Kahneman, 1992), and “DA,” a two-parameter specification of Disappointment Aversion (Gul, 1991), which were studied in FGL. That paper evaluated model restrictiveness in predicting the certainty equivalents for a set of 25 binary lotteries from Bruhin, Fehr-Duda and Epper (2010). Here we are able to evaluate model restrictiveness in predicting certainty equivalents for any lottery, using our general framework.
@@
We define a constrained GP prior $\lambda_{\mathcal{F}}$ on the prediction rule $f$. To sample $f$, we proceed in two steps. First, we draw a monotone increasing function $g(\bar{z},\underline{z},p)$ from a constrained Gaussian process with a Matérn $3/2$ kernel.  [^5] As robustness checks, we (i) replace the Matérn $3/2$ kernel with a squared exponential kernel, and (ii) replace the GP draws with spline basis draws; the qualitative ranking of restrictiveness across models is unchanged. See Online Appendix F.1. Second, we map $g$ to $f$ via the sigmoid transformation $f(\bar{z},\underline{z},p)=\underline{z}+(\bar{z}-\underline{z})\sigma(g(\bar{z},\underline{z},p)),$ ensuring that $f$ takes values in $\underline{z},\bar{z}]$. We use $M=2000$ draws from $\lambda_{\mathcal{F}}$ in our implementation. Random samples from the constrained GP prior satisfy the required monotonicity constraints and exhibit non-flat behavior; see Online Appendix F.1 for illustrative plots. The discrepancy between $f_{1}$ and $f_{2}$ is defined as $L^{2}$-norm of their difference.
@@ 5.2 Multinomial Choice Models
Define the model as the prediction rule $p_{\theta}$ that maps covariates $X_{m}$ in market $m$ to a $J\times 1$ vector of product share $p_{m}(X_{m};\theta)=(p_{1m}(X_{m};\theta),\ldots,p_{Jm}(X_{m};\theta))^{\prime}.$ The baseline model class is a singleton consisting of the uniform share vector $p_{base}=1/(J+1)\cdot\iota$, where $\iota$ is a $J\times 1$ vector. We evaluate the restrictiveness of three widely used discrete-choice specifications: multinomial logit (MNL), nested logit (NL), and mixed logit (MXL), which differ in the substitution patterns they allow: MNL implies independence of irrelevant alternatives, NL allows within-nest correlation, and MXL incorporates random taste heterogeneity. Formal model definitions are given in Appendix F.
@@
In this setting, an eligible set $\mathcal{F}$ is a collection of prediction rules $s$ mapping covariates $X_{m}$ to market-share vectors $s_{m}$. Our specification of the eligible set here is motivated by theoretical work on the flexibility of mixed logit models. Starting from a parametric MXL structure, we consider three variants that differ in which components of utility are allowed to be general functions of product characteristics $x_{jm}$, possibly subject to monotonicity restrictions. Specifically, we consider: (i) an eligible set in which both the common mean utility and individual heterogeneity are allowed to be general functions (“NP Both”), (ii) an eligible set in which the mean utility is allowed to be a general function while individual heterogeneity remains parametric (“NP Mean”), and (iii) an eligible set in which the mean utility remains parametric while individual heterogeneity is allowed to be a general function (“NP Individual”). We introduce an evaluation distribution $\lambda_{\mathcal{F}}$ over each eligible set, defined using (constrained) Gaussian process priors [^6] Robustness checks that vary the kernel choice and the function-draw procedure yield similar results; see Online Appendix F.2., and draw 100 functions from this distribution. Full details are provided in Appendix F.
@@ Results
Table 2 reports restrictiveness for the multinomial choice models under three eligible sets. When both the mean and the individual heterogeneity are modeled nonparametrically (NP Both), all three models remain restrictive, with MNL the most restrictive (0.154) and NL and MXL very similar and less restrictive (0.113 and 0.112). Although mixed logit is theoretically the most flexible, the parametric structure imposed in empirical implementations makes NL and MXL exhibit very similar restrictiveness in our application. Our results also show that the restrictiveness of these models is driven almost entirely by the mean utility component. When the eligible set permits a nonparametric mean utility but keeps individual heterogeneity parametric (NP Mean), restrictiveness is essentially the same as when the eligible set allows both components to be nonparametric (NP Both), and the ranking across models is unchanged. In contrast, when the eligible set keeps the mean utility parametric but allows nonparametric heterogeneity (NP Individual), restrictiveness is near zero for all models, with MXL being the least restrictive. In the “NP Individual” eligible set, flexibility enters only through individual-specific terms and is largely integrated out when forming market shares; as a result, the eligible set is not materially expanded relative to parametric mixed logit. By contrast, allowing a nonparametric mean utility relaxes the functional form of the common utility component and effectively enlarges the eligible set beyond the parametric specification.
Table 2: Restrictiveness of Multinomial Choice Models (No Endogeneity)
Eligible Set Model Restr. SE
NP Both MNL 0.154 0.008
NP Both NL 0.113 0.005
NP Both MXL 0.112 0.005
NP Individual MNL 0.002 0.000
NP Individual NL 0.002 0.000
NP Individual MXL 0.002 0.000
NP Mean MNL 0.157 0.009
NP Mean NL 0.116 0.005
NP Mean MXL 0.120 0.005
Table 3 reports completeness measures, which quantify the fraction of predictable variation in market shares captured by each model relative to a flexible statistical benchmark. [^7] In this setting, a fully nonparametric estimation approach will perfectly fit the observed shares. Using the squared prediction error of market shares as the loss function, we find that models exhibit nearly identical completeness when applied to the cereal data, indicating that the additional parameters in NL and MXL models do not provide meaningful additional predictive power for this dataset.
Table 3: Completeness of Multinomial Choice Models (No Endogeneity)
Model Complete. SE
MNL 0.396 0.039
NL 0.396 0.039
MXL 0.397 0.040

### v3

@@ 4 Related Concepts
This section clarifies the relationship between restrictiveness and related concepts in in econometrics and machine learning.
@@ 4.1 Rademacher Complexity and VC Dimensions
The Rademacher complexity and VC dimension measures implicitly encode a specific discrepancy function that is poorly suited for measuring economic model restrictiveness. This section makes that precise by identifying the discrepancy underlying each capacity measure and tracing the degeneracy result of Ellis and Neff (2025) directly to its properties.
Following Ellis and Neff (2025), suppose that $\mathcal{X}$ is a closed and bounded infinite subset of $\mathbb{R}^{m}$ for some finite $m\in\mathbb{N}$, $X$ is a random vector on $\mathcal{X}$ with distribution $P_{X}$, and the outcome space is $\mathcal{Y}=\{-1,1\}$. [^4] Ellis and Neff allow the outcome $Y$ to take values in [1,1][-1,1]. For expositional purposes, we restrict attention to the canonical binary case, since the most standard definitions of both the Rademacher complexity and the VC dimension are formulated for binary-valued function classes. Let $\bar{\mathcal{F}}_{\sigma}=\{f:\mathcal{X}\to\{-1,1\}\}$ be the set of all deterministic binary classifiers and let $\mathcal{F}_{\Theta}\subseteq\bar{\mathcal{F}}_{\sigma}$ be a model class (for example, a parametric class indexed by $\theta\in\Theta$). Ellis and Neff (2025) take the eligible set to be $\mathcal{F}:=\bar{\mathcal{F}}_{\sigma}$ with some probability measure $\lambda_{\mathcal{F}}$, and the discrepancy to be the correlation-based quantity
@@
As we show in Online Appendix S.2, since $f$ and $g$ are binary-valued, the discrepancy $d_{\text{VC}}(f,g)=\mathbb{P}(f(X)\neq g(X))$ satisfies $d_{\text{Rad}}=2d_{\text{VC}}$, so it induces exactly the same restrictiveness ordering. Thus Rademacher complexity and VC dimension, despite their different motivations, correspond to the same underlying notion of approximation for the purposes of restrictiveness.
Ellis and Neff (2025) establishes that under this discrepancy, all model classes with finite VC dimension satisfy $r=1$ in the limit — they appear maximally restrictive regardless of their actual expressive power. This degeneracy traces directly to the properties of $d_{\text{Rad}}$, which was designed to control worst-case generalization under adversarial labelings, a goal that is orthogonal to comparing the expressive power of economic models.
The choice of discrepancy function is a substantive modeling decision, not a technicality. A discrepancy inherited from an existing capacity measure will reproduce that measure’s asymptotic behavior and limitations, whether or not those properties are desirable for the application at hand. In our applications we use squared $L^{2}$ discrepancies on prediction rules, which measure approximation error in economically interpretable units and do not exhibit this degeneracy. The next subsection applies this same logic to GMM criterion functions, showing that they too are unsuitable as discrepancies and should instead enter the framework as constraints on the eligible set.
@@ 4.2 Discrepancy Associated with GMM Criterion Function
This section explains why GMM criterion functions are not discrepancy functions in the sense required for restrictiveness. It also makes a positive recommendation for how moment conditions should enter the restrictiveness framework.
@@
One might interpret $Q_{\text{GMM}}(\theta)$ as defining a discrepancy between $m_{\theta}$ and the true regression function $g$, and hence to define
@@
unless the instruments are sufficiently rich to identify $g$ pointwise. Thus as shown in the next example $Q_{\text{GMM}}$ does not satisfy $d(f,g)=0\Rightarrow f=g\ \text{(almost surely)}$ which we require of discrepancy functions.
@@ 4.3 Limit of the Average-Case Learning Curve
Proposition 2 shows that under mild assumptions, restrictiveness can be interpreted as the ratio of two average-case approximation errors: one achieved by the best-fitting member of the model class $\mathcal{F}_{\Theta}$, the other by the baseline model class $\mathcal{F}_{\text{base}}$. Low restrictiveness corresponds to a model whose best-fitting members achieve low asymptotic risk across pseudo-truths $f\sim\lambda_{\mathcal{F}}$, while high restrictiveness corresponds to a model that, on average, cannot reduce risk much relative to the baseline. This interpretation is analogous to, but distinct from, the standard notion of a learning curve in machine learning. There, irreducible error arises from the noise term $\epsilon_{i}$ in the data-generating process $Y_{i}=f(X_{i})+\epsilon_{i}$, whereas our setup (22) is noise-free. Restrictiveness can therefore be understood as a normalized, noise-free analog of the limiting average-case learning curve. This perspective reinforces our central point: restrictiveness captures the functional-form flexibility of the model under consideration.
The connection to the Gaussian process (GP) learning-curve literature makes this distinction concrete. In GP regression with squared loss, the average generalization error of the posterior mean (BLUP) takes exactly the form of $L_{n}(\mathcal{F}_{\Theta},\mathscr{A})$ when the average is over both training samples and pseudo-truth draws from the GP prior. In this specialization, Proposition 2 implies that under a noise-free, risk-consistent regime the learning-curve limit equals $\mathbb{E}_{f\sim\lambda_{\mathcal{F}}}[d(\mathcal{F}_{\Theta},f)]$, the pure approximation error. By contrast, the GP learning-curve literature (Le Gratiet and Garnier, 2015) scales the observation-noise variance proportionally to the sample size ($\mathrm{Var}(\epsilon_{i})=n\tau$) so as to produce a nontrivial limit driven by estimation uncertainty rather than model flexibility: even when the GP prior is correctly specified so that $d(\mathcal{F}_{\Theta},f)=0$, the limit is strictly positive. The key difference, therefore, is in what survives in the limit: our noise-free setup isolates the population approximation error (and hence the structural content of the model class), whereas the standard GP limit reflects the fundamental trade-off between learning and noise. See Online Appendix S.3 for further details on the eigenvalue representation and the formal correspondence.
@@ 5 Applications
This section applies our restrictiveness framework to three settings: certainty equivalents, multinomial choice with exogenous product characteristics, and multinomial choice with endogenous prices. In the certainty-equivalent application, the relative importance of parameters mirrors that in FGL, but restrictiveness is uniformly higher when models are evaluated over the full continuum of lotteries. In discrete-choice models without endogeneity, restrictiveness is driven primarily by flexibility in mean utility, and the restrictiveness of commonly used empirical specifications differs meaningfully despite the theoretical generality of mixed logit. When prices are endogenous, moment restrictions substantially increase restrictiveness and alter model rankings. Together, the first two applications illustrate how the framework evaluates restrictiveness for functional models on continuum domains, while the third shows how it extends naturally to structural, semiparametric models with endogeneity.
@@ 5.1 Cumulative Prospect Theory
We revisit the certainty-equivalent setting of Fudenberg, Gao and Liang (2026), using the same models but replacing their finite set of 25 binary lotteries with the full continuum of lotteries. This allows us to isolate the effect of moving from a finite evaluation domain to a functional one, holding the models and discrepancy fixed. We find that restrictiveness is uniformly higher over the continuum, and that the relative contribution of individual parameters is largely preserved, suggesting that FGL’s finite-sample rankings are robust but their absolute levels systematically understate how much structure these models impose.
@@
We define a constrained GP prior $\lambda_{\mathcal{F}}$ on the prediction rule $f$. To sample $f$, we proceed in two steps. First, we draw a monotone increasing function $g(\bar{z},\underline{z},p)$ from a constrained Gaussian process with a Matérn $3/2$ kernel.  [^5] As robustness checks, we (i) replace the Matérn $3/2$ kernel with a squared exponential kernel, and (ii) replace the GP draws with spline basis draws; the qualitative ranking of restrictiveness across models is unchanged. See Online Appendix S.4.1. Second, we map $g$ to $f$ via the sigmoid transformation $f(\bar{z},\underline{z},p)=\underline{z}+(\bar{z}-\underline{z})\sigma(g(\bar{z},\underline{z},p)),$ ensuring that $f$ takes values in $\underline{z},\bar{z}]$. We use $M=2000$ draws from $\lambda_{\mathcal{F}}$ in our implementation. Random samples from the constrained GP prior satisfy the required monotonicity constraints and exhibit non-flat behavior; see Online Appendix S.4.1 for illustrative plots. The discrepancy between $f_{1}$ and $f_{2}$ is defined as $L^{2}$-norm of their difference.
@@ 5.2 Multinomial Choice Models
Define the model as the prediction rule $p_{\theta}$ that maps covariates $X_{m}$ in market $m$ to a $J\times 1$ vector of product shares $p_{m}(X_{m};\theta)=(p_{1m}(X_{m};\theta),\ldots,p_{Jm}(X_{m};\theta))^{\prime}.$ The baseline model class is a singleton consisting of the uniform share vector $p_{base}=1/(J+1)\cdot\iota$, where $\iota$ is a $J\times 1$ vector. We evaluate the restrictiveness of three widely used discrete-choice specifications: multinomial logit (MNL), nested logit (NL), and mixed logit (MXL), which differ in the substitution patterns they allow: MNL implies independence of irrelevant alternatives, NL allows within-nest correlation, and MXL incorporates random taste heterogeneity. Formal model definitions are given in Online Appendix S.4.2.
@@
In this setting, an eligible set $\mathcal{F}$ is a collection of prediction rules $s$ mapping covariates $X_{m}$ to market-share vectors $s_{m}$. Our specification of the eligible set here is motivated by theoretical work on the flexibility of mixed logit models. Starting from a parametric MXL structure, we consider three variants that differ in which components of utility are allowed to be general functions of product characteristics $x_{jm}$, possibly subject to monotonicity restrictions. Specifically, we consider: (i) an eligible set in which both the common mean utility and individual heterogeneity are allowed to be general functions (“NP Both”), (ii) an eligible set in which the mean utility is allowed to be a general function while individual heterogeneity remains parametric (“NP Mean”), and (iii) an eligible set in which the mean utility remains parametric while individual heterogeneity is allowed to be a general function (“NP Individual”). We introduce an evaluation distribution $\lambda_{\mathcal{F}}$ over each eligible set, defined using (constrained) Gaussian process priors [^6] Robustness checks that vary the kernel choice and the function-draw procedure yield similar results; see Online Appendix S.4.2., and draw 1000 functions from this distribution. Full details are provided in Online Appendix S.4.2.
@@ Results
In Panel A of Table 2, the “No Endogeneity” columns report restrictiveness for the multinomial choice models under three eligible sets. When both the mean and the individual heterogeneity are modeled nonparametrically (NP Both), all three models remain restrictive, with MNL the most restrictive (0.154) and NL and MXL very similar and less restrictive (0.113 and 0.112). Although mixed logit is theoretically the most flexible, the parametric structure imposed in empirical implementations makes NL and MXL exhibit very similar restrictiveness in our application. Our results also show that the restrictiveness of these models is driven almost entirely by the mean utility component. When the eligible set permits a nonparametric mean utility but keeps individual heterogeneity parametric (NP Mean), restrictiveness is essentially the same as when the eligible set allows both components to be nonparametric (NP Both), and the ranking across models is unchanged. In contrast, when the eligible set keeps the mean utility parametric but allows nonparametric heterogeneity (NP Individual), restrictiveness is near zero for all models, with MXL being the least restrictive. In the “NP Individual” eligible set, flexibility enters only through individual-specific terms and is largely integrated out when forming market shares; as a result, the eligible set is not materially expanded relative to parametric mixed logit. By contrast, allowing a nonparametric mean utility relaxes the functional form of the common utility component and effectively enlarges the eligible set beyond the parametric specification.
In Panel B of Table 2, the “No Endogeneity” columns report completeness measures, which quantify the fraction of predictable variation in market shares captured by each model relative to a flexible statistical benchmark. [^7] In this setting, a fully nonparametric estimation approach will perfectly fit the observed shares. Using the squared prediction error of market shares as the loss function, we find that models exhibit nearly identical completeness when applied to the cereal data, indicating that the additional parameters in NL and MXL models do not provide meaningful additional predictive power for this dataset.
Table 2: Restrictiveness and Completeness of Multinomial Choice Models
Panel A: Restrictiveness
Eligible Set Model No Endogeneity With Endogeneity Change
Restr. SE Restr. SE $\Delta$
NP Both MNL 0.154 (0.008) 0.745 (0.014) +0.591
NL 0.113 (0.005) 0.745 (0.014) +0.632
MXL 0.112 (0.005) 0.667 (0.022) +0.555
NP Mean MNL 0.156 (0.008) 0.699 (0.010) +0.543
NL 0.115 (0.005) 0.699 (0.010) +0.584
MXL 0.116 (0.005) 0.607 (0.008) +0.491
NP Individual MNL 0.002 (0.000) 0.594 (0.009) +0.592
NL 0.002 (0.000) 0.594 (0.009) +0.592
MXL 0.000 (0.000) 0.548 (0.014) +0.548
Panel B: Completeness
Model No Endogeneity With Endogeneity
Complete. SE Complete. SE
MNL 0.396 (0.039) 0.301 (0.035)
NL 0.396 (0.039) 0.302 (0.035)
MXL 0.397 (0.040) 0.335 (0.040)
Notes: Panel A Column $\Delta$ denotes the difference in restrictiveness between the endogenous and exogenous cases, for each multinomial choice specification and eligible set.